[Brahms-l] Shift report 20050501 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 01 2005 - 08:00:04 EDT
Shifter leader was FR
Shifters were 


 - Beam on from 1:20 am to 7:10 am

 - FS @ 8deg-1723B and MRS @ 45deg-1050A
   3 data runs: 14364, 14365 and 14368. Total of 32.7k trig2 events.
   This brings the total trig2 for this setting to 51k.

   FS @ 8deg-2442B and MRS @ 45deg-1050A
   2 data runs: 14369 and 14370. Total of 4.5k trig2 events

   Pedestal runs: 14362, 14363

 - TPC's and Drift chambers are in park mode

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 1 08:00:30 2005

This archive was generated by hypermail 2.1.8 : Sun May 01 2005 - 08:00:33 EDT